### v2

@@ 4 Systematic detector studies based on simulations
The simulation process illustrated in Section 2 and validated with data in Section 3 can be now used to systematically explore some of the critical parameters of the 3D diamond detectors to seek for possible optimizations. In particular, we want to systematically study the parameters that affect their time resolution in order to get indications for possible design improvements.
@@ 4.1 Effect of electrode resistance and optimization
From the comparison between results obtained with perfectly conductive (i.e. metallic) electrodes and real resistive electrodes, it is evident that the resistance of the graphitized electrodes represents the dominant contribution to the time resolution and is thus the main construction parameter to be optimized.
The most direct effect of high electrode resistance is a proportional increase in the signal time constant, resulting in a longer rise time (). Since the total collected charge, and hence the integral of the signal, remains constant, a longer rise time leads to a reduction in the signal amplitude (). Assuming a constant noise level, the signal-to-noise ratio (SNR) decreases accordingly. The rise time and SNR determine the jitter contribution to the time resolution, given by:
where is the standard deviation of the noise distribution. Thus, an increased electrode resistance directly results in a larger jitter contribution. From Equation 4.1 we can estimate what would be the minimum jitter contribution, with perfectly conductive electrodes. Convolving a very short current spike with the KU board transfer function, and using the measured electronic noise (5 mV RMS) for , we obtain a minimum jitter contribution of about 11 ps. Taking into account also the uncertainty caused by variation in the drift times of charge carriers across a half cell, we estimate a minimum jitter contribution to the time resolution of approximately 15 ps.
With the simulation we have studied in a systematic way the effect of the electrode resistance on the time resolution. The results of this study are reported in Figure 15. This has been achieved by changing the resistance in the discrete impedance model and by scaling the time dependence of the weighting potential (see Section 2.3).
While showing a significant discrepancy at large resistances (due to the over-simplification of the lumped-element model) the two approaches converge to the same asymptotic time resolution below reasonably achievable resistance values of a few tens of k. The asymptotic minimum time resolution is given by the sum in quadrature of the jitter contribution described above and the contribution of the inhomogeneities of the electric field due to the sensor geometry that will be discussed in the next section.
@@ 4.2 Optimization of the geometry
Using the spectral method, modifying the sensor geometry and recomputing time-dependent weighting potential and electrostatic field maps is relatively fast and straightforward. The generated maps can then be fed to Garfield++ without further modifications to the simulation, signal processing and data analysis flow.
To test different geometries we have considered three configurations alternative to the geometry discussed in the previous sections:
parallelepiped electrodes, where all the electrodes are parallelepiped with square section and side 11 m (represented in Figure 16a);
trench electrodes, emulating the geometry of the TimeSPOT silicon sensors [44]. This configuration is most probably unrealistic for diamond without significant breakthrough in the graphitization, due to fabrication limitations (trenching in diamond is indeed prone to induce cracks in the material), but is reported here as an extreme limit to what is achievable by the sole modification of the sensor geometry (represented in Figure 16c);
@@
Trench electrodes |
The resolutions obtained from these simulation runs, assuming either perfectly conducting or graphitized electrodes, are reported in Table 1. The results with metallic electrodes provide an estimate of the contribution of field non-uniformities to the overall time resolution, whereas the results with a realistic resistivity represent a prediction of the sensor’s timing performance achievable with state-of-the-art graphitization technologies.
@@
Metallic trench electrodes, extensively studied by the TimeSPOT collaboration [44], are expected to provide the best performance in terms of field uniformity. This is confirmed by the simulation which estimates a time resolution close to the best theoretically achievable value given the assumed electronics and electronic noise. The residual difference is largely due to the loss of available charge carriers due to the thicker electrodes. Also, the larger thickness reduces the resistance of realistic graphitic electrodes improving the signal propagation characteristics as demonstrated by the significantly better time resolution. However, even with this extreme geometry, the electrode resistance remains the dominant contribution to the sensor timing performance.
@@ 4.3 Time resolution and the effect of tilting angles
This additional degradation has been traced back to Landau fluctuations in the ionization energy deposits along the particle trajectory combined with a variable signal propagation delay that depends on the position at which the signal is induced along the electrode [44]. Indeed, Landau fluctuations in the ionization deposits cause significant variability in the distribution of carrier densities along the electrode length, leading to a variable mixture of signal components subject to different delays. This variability in the charge distribution is then reflected in fluctuations of the signal’s rising edge and, ultimately, in the time stamp assigned to the waveform by the signal processing algorithm.
@@
The variability in the distribution of the charge induced along the electrode length can be enhanced by tilting the sensor axis with respect to the impinging particles. In silicon 3D sensors, where the timing resolution is mainly limited by inhomogeneities of the electrostatic field, tilting the sensor is generally beneficial. This is because it reduces the likelihood that the trajectory of an incoming particle lies entirely within either a high- or low-field region, thereby mitigating discrepancies between tracks, as most will traverse both types of regions.
In the case of 3D diamond sensors, however, tilting the sensor causes particles to traverse high- and low-field regions at different depths, increasing the variability of propagation delays of induced signals along the electrode and, consequently, of the resulting leading edges.
The simulation developed in this work provides the opportunity to investigate the time resolution as a function of sensor inclination. This study is important for identifying the optimal operating angle that maximizes timing performance by balancing the competing effects of field inhomogeneities and fluctuations in charge deposition along the particle track.
@@
This study highlights an effect of column resistance that extends beyond jitter and remains independent of electronic noise. Furthermore, it suggests that aligning diamond sensor electrodes with the expected trajectory of impinging particles may enhance timing resolution. In contrast, silicon 3D sensors exhibit improved performance when electrodes have an inclination of 10°or more with respect to the incident particles. This configuration, however, promotes charge sharing between adjacent pixels, which can potentially degrade overall time resolution [47].
@@ 5 Conclusion
In this work, we have provided a detailed analysis of the contributions to the timing uncertainty in 3D diamond detectors with graphitic electrodes. To this end we have implemented a full simulation workflow exploiting a generalized version of the Ramo-Shockley theorem, developed for conductive media to account for signal propagation effects within the impedance network of a realistic sensor. Specifically, the simulation combines the modeling of ionization charge deposits performed with Heed and the computation of carrier trajectories implemented in Garfield++, with time-dependent weighting potentials obtained by numerically solving the Maxwell equations in the quasi-static limit either via Finite Element Methods using COMSOL MultiPhysics® or through a custom solver leveraging spectral methods. The results obtained with these two alternative methods are reasonably consistent and qualitatively agree with expectations from a simplified model representing the sensor as a discrete impedance network.
@@
By varying the simulated electrode resistivity we observed that already at 0.3 the contribution from the signal propagation becomes subdominant relative to the combined effects of jitter and field inhomogeneities.

### v3

@@ 4 Systematic detector studies based on simulations
The simulation process illustrated in Section 2 and compared with real data in Section 3 can be now used to systematically explore some of the critical parameters of the 3D diamond detectors to seek for possible optimizations. In particular, we want to systematically study the parameters that affect their time resolution in order to get indications for possible design improvements.
@@ 4.1 Effect of electrode resistance and optimization
From the comparison between results obtained with metallic electrodes and real resistive electrodes, it is evident that the resistance of the graphitized electrodes represents the dominant contribution to the time resolution and is thus the main construction parameter to be optimized.
The most direct effect of high electrode resistance is a proportional increase in the signal time constant, resulting in a longer rise time (), and to a lesser extent, fall time (). Since the total collected charge, and hence the integral of the signal, remains constant, longer rise (and fall) time leads to a reduction in the signal amplitude (). Assuming a constant noise level, the signal-to-noise ratio (SNR) decreases accordingly. The rise time and SNR determine the jitter contribution to the time resolution, given by:
where is the standard deviation of the noise distribution. Thus, an increased electrode resistance directly results in a larger jitter contribution. From Equation 4.1 we can estimate what would be the minimum jitter contribution, with perfectly conductive electrodes. Convolving a very short current spike with the KU board transfer function, resulting in a waveform with and , and using the measured electronic noise (5 mV RMS) for , we obtain a minimum jitter contribution of about 11 ps. Taking into account the uncertainty caused by variation in the drift times of charge carriers across a half cell, we estimate a minimum jitter contribution to the time resolution of approximately 15 ps.
With the simulation, we have studied in a systematic way the effect of the electrode resistance on the time resolution. The results of this study are reported in Figure 15. This has been achieved by changing the resistance in the discrete impedance model and by scaling the time dependence of the weighting potential (see Section 2.4).
While showing a significant discrepancy at large resistances (primarily due to the over-simplification of , especially in the early rise of the signal which is more relevant for time resolution) the two approaches converge to the same asymptotic time resolution below reasonably achievable resistance values of a few tens of k. The asymptotic minimum time resolution is given by the sum in quadrature of the jitter contribution described above and the contribution of the inhomogeneities of the electric field due to the sensor geometry that will be discussed in the next section. The agreement at low resistance is not surprising, as process 3, the only one treated differently in and , becomes progressively less relevant in that regime.
@@ 4.2 Optimization of the geometry
Using the spectral method (), modifying the sensor geometry and recomputing time-dependent weighting potential and electrostatic field maps is relatively fast and straightforward. The generated maps can then be fed to Garfield++ without further modifications to the simulation, signal processing and data analysis flow.
To test different geometries we have considered three configurations alternative to the geometry discussed in the previous sections and represented in Figure 16a:
parallelepiped electrodes, where all the electrodes are parallelepiped with square section and side 11 m (represented in Figure 16b), this geometry essentially corresponds to the cylindrical configuration used in the COMSOL MultiPhysics® field maps;
tilted electrodes, where the cylindrical polarization electrodes, with 12 m diameter, are tilted to favor field uniformity as proposed in Ref. [HuazhenLi:2024vci] (represented in Figure 16c);
trench electrodes, emulating the geometry of the TimeSPOT silicon sensors [Anderlini:2020ffm]. This configuration is most probably unrealistic for diamond without significant breakthrough in the graphitization, due to fabrication limitations (trenching in diamond is indeed prone to induce cracks in the material), but is reported here as an extreme limit to what is achievable by the sole modification of the sensor geometry (represented in Figure 16d);
@@
| Trench electrodes | ||
The resolutions obtained from these simulation runs, assuming either perfectly conducting or graphitic electrodes, are reported in Table 1. The results with metallic electrodes provide an estimate of the contribution of field non-uniformities to the overall time resolution, whereas the results with a realistic resistivity represent a prediction of the sensor’s timing performance achievable with state-of-the-art graphitization technologies.
@@
Metallic trench electrodes, extensively studied by the TimeSPOT collaboration [Anderlini:2020ffm], are expected to provide the best performance in terms of field uniformity. This is confirmed by the simulation which estimates a time resolution close to the best theoretically achievable value given the assumed electronics and electronic noise. The residual difference is largely due to the loss of available charge carriers due to the thicker electrodes. Also, the larger thickness reduces the resistance of realistic graphitic electrodes improving the signal propagation characteristics as demonstrated by the significantly better time resolution. However, even with this extreme geometry, the electrode resistance remains the dominant contribution to the sensor timing performance.
@@ 4.3 Time resolution and the effect of tilting angles
This additional degradation has been traced back to Landau fluctuations in the ionization energy deposits along the particle trajectory combined with a variable signal propagation delay that depends on the position at which the signal is induced along the readout electrode [Anderlini:2020ffm]. Indeed, Landau fluctuations in the ionization deposits cause significant variability in the distribution of carrier densities along the electrode length, leading to a variable mixture of signal components subject to different delays. This variability in the charge distribution is then reflected in fluctuations of the rising edge of the signal and, ultimately, in the time stamp assigned to the waveform by the signal processing algorithm.
@@
The distribution of the charge induced along the electrode length is influenced by the tilt of the incoming particle relative to the readout electrodes. In silicon 3D sensors, where the timing resolution is mainly limited by inhomogeneities of the electrostatic field, tilting the sensor is generally beneficial. This is because it reduces the probability that the trajectory of an incoming particle lies entirely within either a high- or low-field region, thereby mitigating discrepancies between tracks, as most will traverse both types of regions.
In contrast, for 3D diamond sensors with resistive electrodes tilting the sensor causes particles to traverse high- and low-field regions at different depths, increasing the variability of propagation delays of induced signals along the electrode. Consequently, the leading edges of the signals become more variable, and the time resolution deteriorates as the inclination angle increases.
The simulation developed in this work provides the opportunity to investigate the time resolution as a function of sensor inclination. In particular, it allows to disentangle two competing effects: one beneficial, related to the exploration of regions with different electric fields during the particle traversal, and one detrimental, arising from the tilt-induced variability of delayed signal components along the electrode. This study is important for identifying the optimal operating angle that maximizes timing performance by balancing the competing effects of field inhomogeneities and fluctuations in charge deposition along the particle track.
@@
This study highlights an effect of column resistance that extends beyond jitter and remains independent of electronic noise. Furthermore, it suggests that aligning diamond sensor electrodes with the expected trajectory of impinging particles may enhance timing resolution. In contrast, silicon 3D sensors exhibit improved performance when electrodes have an inclination of 10° or more with respect to the incident particles. This configuration, however, promotes charge sharing between adjacent pixels, which can potentially degrade overall time resolution [Timespot:2023].
@@ 5 Conclusion
In this work, we have provided a detailed analysis of the contributions to the timing uncertainty in 3D diamond detectors with graphitic electrodes. To this end we have implemented a full simulation workflow exploiting a generalized version of the Ramo-Shockley theorem, developed for conductive media to account for signal propagation effects within the impedance network of a realistic sensor. Specifically, the simulation combines the modeling of ionization charge deposits performed with Heed and the computation of carrier trajectories implemented in Garfield++, with time-dependent weighting potentials obtained by numerically solving the Maxwell equations in the quasi-static limit either via FEMs using COMSOL MultiPhysics® () or through a custom solver leveraging spectral methods (). The results obtained with these two alternative methods are reasonably consistent and qualitatively agree with expectations from a simplified model representing the sensor as a discrete impedance network ().
@@
By varying the simulated electrode resistivity to assess the potential benefits of improved conductivity, we observed that already at 0.3 and below the contribution from the signal propagation becomes subdominant relative to the combined effects of jitter and field inhomogeneities.